The points A,B,C,D,E are marked on the circumference of a circle in clockwise direction such that angle ABC = 130° and angle CDE = 110° . The measure of angle ACE in degrees is A. 50°B. 60°C. 70°D. 80°

Grade:11

2 Answers

Arun
6 years ago
Use sum of the arcs is always 360° and angle incribed in a arc is always half of the arc inscribed in it.
you will get answer as 60°.
Charan Raju
6 years ago
AngAEC +AngABC =180degrees (sum of Opposite angles in a cyclic quadrilateral is 180degrees)=> AngAEC = 180-AngABC=180-130 = 50degreesAngCAE+AngCDE = 180degrees (same reason)AngCAE = 180-AngCDE= 180-110=70degreesAngCAE+AngAEC+AngACE=180degress (sum of angles in a triangle is 180)=> AngACE = 180-(50+70)=180-120 = 60degrees